OpenStudy (anonymous):

choose 2 fractions between 2/7 & 3/7

OpenStudy (lilg132):

3/8

OpenStudy (lilg132):

dont forget to click good answer :)

OpenStudy (anonymous):

I dunt think this is a good answer. CHOOSE 2 FRACTIONS, NOT 1.

OpenStudy (lilg132):

oh sorry i misread the question

OpenStudy (anonymous):

ITS OK

OpenStudy (lilg132):

2/7 = 0.2857142857142857 3/7 = 0.42857142857142855 3/8 = 0.375 1/3 = 0.3333333333333333 both between the two fractions :)

OpenStudy (anonymous):

choose one of this answer 7/7+1; 1/7 21/70; 26/63

OpenStudy (anonymous):

or 3*2/7*7; 3+2/7*7 3+2/7+7; 2/7+1

OpenStudy (lilg132):

you havent clicked good answer yet

OpenStudy (anonymous):

cuz, it's not a gud answer

OpenStudy (anonymous):

sorry:((

OpenStudy (lilg132):

you said choose 2 fractions and i gave you 2 fractions

OpenStudy (anonymous):

yup, but its not a gud answer. choose one of this answer 7/7+1; 1/7 21/70; 26/63 or 3*2/7*7; 3+2/7*7 3+2/7+7; 2/7+1

OpenStudy (lilg132):

21/70; 26/63

OpenStudy (lilg132):

still not going to click good answer

OpenStudy (anonymous):

\[\left\{\frac{2}{7},\frac{3}{7}\right\}=\{0.285714,0.428571\} \]\[\left\{\frac{2}{7},\frac{3}{10},\frac{4}{10},\frac{3}{7}\right\}=\{0.285714,0.3,0.4,0.428571\} \]